molly is 4 years younger than heidi.in 3 years,heidi will be twice as old as molly.how old are they now ?
Found 2 solutions by oberobic, richwmiller:
Answer by oberobic(2304)

m = Molly's age now.
h = Heidi's age now.
Molly is 4 years younger than Heidi.
So, if you subtract 4 from Heidi's age, you will know Molly's age.
m = h-4
.
"In three years..."
(m+3) will be Molly's age then.
(h+3) will be Heidi's age then.
(h+3) = 2(m+3) :: given.
.
Subsitute what we know about m, which is it equal h-4.
.
(h+3) = 2(m+3)
(h+3) = 2m + 6
substitute m=h-4
(h+3) = 2(h-4) +6
h +3 = 2(h -4) +6
h +3 = 2h -8 +6
h +3 = 2h -2
h = 2h -8 +3
-h = -5
h = 5
Heidi is 5 years old now.
.
m = h-4
m = 5-4
m = 1
Molly is 1 years old now.
.
Check this answer by testing how old they will be in 3 years?
Molly will be 8.
Heidi will be 4.
Will Molly be twice as old as Heidi?
Yes.
.
Done.

Answer by richwmiller(17219)

m=h-4
h=m+4
h+3=2m
h=2m-3
2m-3=m+4
m=7
h=11